OpenStudy (anonymous):

Widget Wonders produces widgets. They have found that the cost, c(x), of making x widgets is a quadratic function in terms of x. The company also discovered that it costs $29 to produce 2 widgets, $115 to produce 4 widgets, and $757 to produce 10 widgets. Complete the function that represents the cost, c(x), to produce x widgets.

OpenStudy (thomas5267):

Let \(c(x)=ax^2+bx+c\). Put the three sets of value in and solve it.

OpenStudy (anonymous):

So, c(x)=29x^2+115x+757?

OpenStudy (thomas5267):

No. The first set of conditions is \(c(2)=27\). Can you fnd the other ones?

OpenStudy (anonymous):

c(4)=115?

OpenStudy (thomas5267):

And the last one?

OpenStudy (anonymous):

c(10)=757

OpenStudy (thomas5267):

Okay, put that back into \(c(x)=dx^2+ex+f\). I changed a, b and c to d, e and f because the function is c and the constant is c too. That is a bit confusing. The first equation would be \(2^2d+2e+f=29\). Just create the three equations and solve it.

OpenStudy (anonymous):

So the second one would be 4x^2+4x+c=115?

OpenStudy (thomas5267):

No. You know that c(4)=115 and \(c(x)=dx^2+ex+f\). You replace all x in c(x) with 4 as you want c(4). That is: \[ c(4)=d4^2+e4+f=115\\ =16d+4e+f=115\]

OpenStudy (thomas5267):

@Callisto I am now at Sogo because of the sale and I don't think it is a good idea to do OpenStudy while walking in a crowded area so please help!

OpenStudy (callisto):

Aiya! Sorry for the belated reply. As Thomas has mentioned, since the function is quadratic in x, we have c(x) = ax^2 + bx + c Also, we are given when 2 widgets are produced, the cost is $29, i.e. c(2) = 29. When 4 widgets are produced, the cost is $115,i.e. c(4) = 115 and when 10 widgets are produced, the cost is $757, i.e. c(10)=757. So, we have the system: 29 = (2^2) a + 2b + c 115 = (4^2) a + 4b + c 757 = (10 ^2 ) a + 10b + c Now, you need to simply the equations and solve the system.

OpenStudy (anonymous):

thank you guys so much!!! @callisto @thomas5267